5(x+2)-7=5x+3
Solution with steps please?

Answer by Ave(106)

5(x+2)-7=5x+3 get rid of parenthesis by multiplying each term inside by 5
5x+10-7=5x+3 combine like terms
5x+3=5x+3
Always True or "all x" just means that any number for x works.